# Skald: Learning-Based Shot Assembly for Coherent Multi-Shot Video Creation

### v1

@@ Abstract
We present Skald, a multi-shot video assembly method that constructs coherent video sequences from candidate shots with minimal reliance on text. Central to our approach is the Learned Clip Assembly (LCA) score, a learning-based metric that measures temporal and semantic relationships between shots to quantify narrative coherence. We tackle the exponential complexity of combining multiple shots with an efficient beam-search algorithm guided by the LCA score. To train our model effectively with limited human annotations, we propose two tasks for the LCA encoder: Shot Coherence Learning, which uses contrastive learning to distinguish coherent and incoherent sequences, and Feature Regression, which converts these learned representations into a real-valued coherence score. We develop two variants: a base Skald model that relies solely on visual coherence and Skald-text, which integrates auxiliary text information when available. Experiments on the VSPD and our curated MSV3C datasets show that Skald achieves an improvement up to 48.6% in IoU and a 43% speedup over the state-of-the-art methods. A user study further validates our approach, with 45% of participants favoring Skald-assembled videos, compared to 22% preferring text-based assembly methods.
@@ 1 Introduction
Figure 2: Challenges in multi-shot video assembly. Text-based retrieval selects relevant shots independently but may create incoherent sequences (e.g., a snow scene in summer footage).
@@ 2.1 Text-guided Shot Assembly
Several works have addressed the problem of shot assembly through text-to-video retrieval [34, 25, 16, 33]. Given text descriptions, either at the shot level or as a full video synopsis, these approaches encode both text and video modalities into a shared representation space, finding the closest matching embeddings. Unlike the traditional text-video retrieval problem, this line of work also emphasizes intra-shot cohesion, ensuring that the retrieved shots are coherent with each other. For instance, Yang et al. [34] propose transformer-based encoders that take both text and video tokens as input, training the encoder to distinguish reference videos from shuffled or randomly inserted shot sequences using cross-entropy loss. Xiong et al. [33] introduce an adaptive encoder module that improves video-to-video coherence by updating text embeddings based on visual cues from preceding frames, enhancing coherence. Lu et al. [16] further improve consistency by aligning text and video representations at both local (shot-to-text) and global (story-level) scales. Despite their effectiveness, these approaches rely on expensive shot-level text annotations and custom datasets, limiting adaptability to new video domains and constraining real-world applicability. In contrast, we directly score multi-shot video sequences for coherence, guiding assembly without detailed text labels.
@@ 2.2 Video Quality Assessment (VQA)
Extensive work has been done on developing quality metrics for videos to assess their technical quality and aesthetics. Most approaches involve training neural networks to align with human preferences. Li et al[11] proposed VSFA, which incorporates content-aware features from pre-trained CNNs and GRU-based temporal modeling, along with a subjectively-inspired temporal pooling layer human-centric quality perception. Wu et al. [27] developed FAST-VQA, an efficient end-to-end VQA method that uses fragment sampling to reduce computational costs while maintaining accuracy. In a more holistic approach, Wu et al[28] explored evaluating both technical and aesthetic aspects of video quality. They created the DIVIDE-3k dataset—offering diverse human opinions on video quality—and proposed DOVER, a model addressing both technical and aesthetic facets. Building on this, Wu et al[29] introduced Q-Align, leveraging large multi-modal models for visual quality assessment, demonstrating robust generalization to various visual tasks. However, existing VQA methods typically demand large-scale human preference data and assess individual shot quality rather than multi-shot coherence. As a result, well-shot yet thematically disjoint clips can receive high VQA scores despite lacking narrative structure. To address this, we propose the Learned Clip Assembly (LCA) score, a novel metric for evaluating coherence, logical flow, and overall assembly quality. Our two-step training—contrastive pre-training followed by minimum human supervision—significantly reduces annotation needs while effectively capturing multi-shot coherence.
@@ 3.4 Skald: Multi-Shot Video Assembly
The time complexity of this beam search is $\mathcal{O}\bigl{(}k^{3}+(n-3)\,m\,k\bigr{)},$ where $k^{3}$ arises from the initial exhaustive search of the first three shots, and $(n-3)\,m\,k$ covers evaluating $k$ extensions for each of the $m$ beam elements for the remaining $(n-3)$ positions. This is substantially more efficient than the naive $\mathcal{O}(k^{n})$ search when $m\ll k^{\,n-3}$, making it feasible for real-world multi-shot video assembly tasks.
@@ 5 Conclusion
We present Skald, a multi-shot video assembly framework designed to create coherent video sequences from candidate shots with minimal reliance on video captions, while optionally incorporating textual cues through its Skald-text variant. At the core of Skald is the Learned Clip Assembly (LCA) score, a metric trained to quantify narrative coherence by modeling temporal and semantic relationships between shots. Our two-stage approach—contrastive pretraining for shot coherence followed by supervised regression—enables efficient training even under limited annotation budgets. Leveraging beam search enhanced with sequence-length decorrelation, Skald efficiently navigates the combinatorial shot space without exhaustive enumeration. Experiments on the MSV3C and VSPD datasets demonstrate substantial improvements over prior methods in both video quality and computational efficiency, validated by a user study indicating clear subjective preference. As future work, we aim to apply Skald to longer-form contents, enabling automatic assembly of complex narratives in scenarios such as short films or documentary-style storytelling.

### v2

@@ Abstract
We present Skald, a multi-shot video assembly method that constructs coherent video sequences from candidate shots with minimal reliance on text. Central to our approach is the Learned Clip Assembly (LCA) score, a learning-based metric that measures temporal and semantic relationships between shots to quantify narrative coherence. We tackle the exponential complexity of combining multiple shots with an efficient beam-search algorithm guided by the LCA score. To train our model effectively with limited human annotations, we propose two tasks for the LCA encoder: Shot Coherence Learning, which uses contrastive learning to distinguish coherent and incoherent sequences, and Feature Regression, which converts these learned representations into a real-valued coherence score. We develop two variants: a base Skald model that relies solely on visual coherence and Skald-text, which integrates auxiliary text information when available. Experiments on the VSPD and our curated MSV3C datasets show that Skald achieves an improvement up to 48.6% in IoU and a 43% speedup over the state-of-the-art methods. A user study further validates our approach, with 45% of participants favoring Skald-assembled videos, compared to 22% preferring text-based assembly methods. Project URL: https://schaterji.io/publications/2025/SKALD
@@ 1 Introduction
Figure 2: Challenges in multi-shot video assembly. Text-based retrieval selects relevant shots independently but may create incoherent sequences (e.g., a snow scene amid summer footage).
@@ 2.1 Text-guided Shot Assembly
Several works have addressed the problem of shot assembly through text-to-video retrieval [34, 25, 16, 33]. Given text descriptions, either at the shot level or as a full video synopsis, these approaches encode both text and video modalities into a shared representation space, finding the closest matching embeddings. Unlike the traditional text-video retrieval problem, this line of work also emphasizes intra-shot cohesion, ensuring that the retrieved shots are coherent with each other. For instance, Yang et al[34] propose transformer-based encoders that take both text and video tokens as input, training the encoder to distinguish reference videos from shuffled or randomly inserted shot sequences using cross-entropy loss. Xiong et al[33] introduce an adaptive encoder module that improves video-to-video coherence by updating text embeddings based on visual cues from preceding frames, enhancing coherence. Lu et al[16] further improve consistency by aligning text and video representations at both local (shot-to-text) and global (story-level) scales. Despite their effectiveness, these approaches rely on expensive shot-level text annotations and custom datasets, limiting adaptability to new video domains and constraining real-world applicability. In contrast, we directly score multi-shot video sequences for coherence, guiding assembly without detailed text labels.
@@ 2.2 Video Quality Assessment (VQA)
Extensive work has been done on developing quality metrics for videos to assess their technical quality and aesthetics. Most approaches involve training neural networks to align with human preferences. Li et al[11] proposed VSFA, which incorporates content-aware features from pre-trained CNNs and GRU-based temporal modeling, along with a subjectively-inspired temporal pooling layer human-centric quality perception. Wu et al[27] developed FAST-VQA, an efficient end-to-end VQA method that uses fragment sampling to reduce computational costs while maintaining accuracy. In a more holistic approach, Wu et al[28] explored evaluating both technical and aesthetic aspects of video quality. They created the DIVIDE-3k dataset—offering diverse human opinions on video quality—and proposed DOVER, a model addressing both technical and aesthetic facets. Building on this, Wu et al[29] introduced Q-Align, leveraging large multi-modal models for visual quality assessment, demonstrating robust generalization to various visual tasks. However, existing VQA methods typically demand large-scale human preference data and assess individual shot quality rather than multi-shot coherence. As a result, well-shot yet thematically disjoint clips can receive high VQA scores despite lacking narrative structure. To address this, we propose the Learned Clip Assembly (LCA) score, a novel metric for evaluating coherence, logical flow, and overall assembly quality. Our two-step training—contrastive pre-training followed by minimum human supervision—significantly reduces annotation needs while effectively capturing multi-shot coherence.
@@ 3.4 Skald: Multi-Shot Video Assembly
The time complexity of this beam search is $\mathcal{O}\bigl{(}k^{3}+(n-3)\,m\,k\bigr{)},$ where $k^{3}$ arises from the initial exhaustive search of the first three shots, and (n3)mk(n-3)\,m\,k covers evaluating $k$ extensions for each of the $m$ beam elements for the remaining (n3)(n-3) positions. This is substantially more efficient than the naive $\mathcal{O}(k^{n})$ search when $m\ll k^{\,n-3}$, making it feasible for real-world multi-shot video assembly tasks.
@@ 5 Conclusion
We present Skald, a multi-shot video assembly framework designed to create coherent video sequences from candidate shots with minimal reliance on video captions, while optionally incorporating textual cues through its Skald-text variant. At the core of Skald is the Learned Clip Assembly (LCA) score, a metric trained to quantify narrative coherence by modeling temporal and semantic relationships between shots. Our contrastive pretraining for shot coherence followed by supervised regression enables efficient training even under limited annotation budgets. Leveraging beam search enhanced with sequence-length decorrelation, Skald efficiently navigates the combinatorial shot space without exhaustive enumeration. Experiments on the MSV3C and VSPD datasets demonstrate substantial improvements over prior methods in both video quality and computational efficiency, validated by a user study indicating clear subjective preference. We aim to apply Skald to longer-form contents as future work.
Acknowledgement
This material is based in part upon work supported by the National Science Foundation under NSF Grant Numbers CNS-2146449 (NSF-CAREER) and CNS-2333487 (NSF-FRONTIER) and gift funding from Adobe Research. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors.
Supplementary Material
Figure 7: Qualitative analysis shows that standard text-based retrieval methods readily match individual pizza-making steps but struggle to maintain cohesive transitions between them. By contrast, our approach consistently assembles visually coherent sequences that accurately follow each stage of dough preparation.
Figure 8: Qualitative analysis indicates that basic text-based methods identify the individual cleaning tasks but often lack seamless transitions between them. In contrast, our approach constructs visually coherent sequences that reflect each stage of kitchen organization.
Figure 9: Qualitative analysis indicates that basic text-based methods identify the individual cleaning tasks but often lack seamless transitions between them. Nonetheless, our method ensures smooth transition between shots for a promotion video on environment awareness.
6 More Details on Data Preparation
MSV3C training set.
We collected videos from the V3C1 dataset [22], comprising $28,450$ videos of 252-5 minutes duration. While V3C1 provides scene segmentation, we found its scene cuts did not accurately align with shot transitions. Therefore, we used TransNetV2 [23] to split videos into individual shots. We filtered out shots shorter than 2 seconds, static frames, and shots containing excessive text overlays. To ensure video diversity, we computed the average CLIP embedding for each shot and analyzed variance across videos. We excluded the lowest variance quartile, typically consisting of interview-style videos with limited visual diversity. For each remaining video, we randomly sampled between 3 and 12 consecutive shots, forming representative multi-shot sequences suitable for training and evaluation. The resulting dataset encompasses 5,648 carefully curated videos, totaling over 12,000 shots.
MSV3C testing set.
We divided the MSV3C dataset into training and testing sets with a 9:1 ratio. To simulate real-world video editing scenarios, each shot in the test set was manually annotated with a concise single-sentence scene description. These descriptions served as textual queries to retrieve five visually similar candidate shots per original shot from our internal stock footage collection. This design closely reflects practical shot assembly workflows, where the objective is identifying the optimal coherent sequence among visually and semantically related alternatives.
7 User Study
We conducted a comprehensive user study to rigorously assess and compare the video assembly quality between CLIP4Clip [17] and Skald. The primary objective was evaluating these methods’ ability to generate coherent and engaging visual narratives without relying on audio elements.
Our study employed a side-by-side comparison format. Participants simultaneously viewed two multi-shot videos sharing identical thematic content. Each video explicitly excluded voice-overs and music, enabling unbiased evaluation of visual storytelling proficiency. Participants independently rated the quality of each video on a standardized five-point scale from “Very Poor” to “Excellent,” ensuring precise quantification of their qualitative judgments. Participants were provided with ample time—up to 30 minutes—although task completion typically required less than half that duration.
To maintain data reliability, “golden set” video pairs with predetermined ratings were included as quality controls. Submissions failing these controls were excluded. The study involved 3 distinct rating sessions covering 10 unique video pairs, with each pair evaluated multiple times by different participants. This robust redundancy mitigated subjective bias, ensuring comprehensive evaluation.
8 More Qualitative Results
The qualitative analysis presented in Fig. 7Fig. 8, and Fig. 9 underscores that Skald consistently outperforms conventional text-based methods in preserving visual continuity and narrative coherence. As illustrated in Fig. 7, traditional text-based retrieval successfully identifies individual pizza-making steps but frequently introduces visual discontinuities such as inconsistent lighting conditions or contextually irrelevant scenes. Conversely, Skald reliably assembles sequences that flow logically through each cooking stage, sustaining thematic unity. Similarly, in Fig. 8, standard retrieval methods often correctly pinpoint discrete kitchen-cleaning tasks yet fail to ensure seamless visual and contextual transitions. Skald excels by selecting visually coherent and contextually consistent shots, effectively conveying a natural narrative progression. Finally, the environmental awareness promotion in Fig. 9 exemplifies Skald’s capacity to preserve smooth transitions across visually distinct but semantically connected scenes, significantly enhancing viewer engagement through sustained narrative coherence.
9 Video Demo
We provide three example videos (0001.mp4, 0002.mp4, 0003.mp4) assembled with Skald. The detailed themes and descriptions corresponding to each scene within these videos are comprehensively documented in the accompanying file video demo.rtf. These examples demonstrate Skald ’s practical capability in generating high-quality, thematically cohesive multi-shot videos across diverse scenarios.